# Task Aware Dreamer for Task Generalization in Reinforcement Learning

### v3

@@ Abstract
A long-standing goal of reinforcement learning is to acquire agents that can learn on training tasks and generalize well on unseen tasks that may share a similar dynamic but with different reward functions. The ability to generalize across tasks is important as it determines an agent’s adaptability to real-world scenarios where reward mechanisms might vary. In this work, we first show that training a general world model can utilize similar structures in these tasks and help train more generalizable agents. Extending world models into the task generalization setting, we introduce a novel method named Task Aware Dreamer (TAD), which integrates reward-informed features to identify consistent latent characteristics across tasks. Within TAD, we compute the variational lower bound of sample data log-likelihood, which introduces a new term designed to differentiate tasks using their states, as the optimization objective of our reward-informed world models. To demonstrate the advantages of the reward-informed policy in TAD, we introduce a new metric called Task Distribution Relevance (TDR) which quantitatively measures the relevance of different tasks. For tasks exhibiting a high TDR, i.e., the tasks differ significantly, we illustrate that Markovian policies struggle to distinguish them, thus it is necessary to utilize reward-informed policies in TAD. Extensive experiments in both image-based and state-based tasks show that TAD can significantly improve the performance of handling different tasks simultaneously, especially for those with high TDR, and display a strong generalization ability to unseen tasks.
(a) Overview of this work
(b) Probabilistic graphical model designs
Figure 1: (a) An overview. Given a task distribution, we train the agent in training tasks and hope it to zero-shot generalize to testing tasks. For improving the generalization, we propose TAD, which utilizes $\Pi_{3}$ to encode all historical information for inferring the current task and novel reward-informed world models for capturing invariant latent features. (b) Probabilistic graphical model designs for the single-task setting (left) and the task-distribution setting (right). The latter inspires the design of reward-informed world models. Solid and dashed lines represent the generative process and the inference model, respectively.
@@ 1 Introduction
Deep Reinforcement Learning (DRL) has demonstrated significant advancements in diverse fields Mnih et al. (2016); Hafner et al. (2019a) , and a key factor in these accomplishments is an agent’s proficiency in assimilating lessons from dedicated training tasks. However, when the objective transitions from optimizing performance in specific tasks to achieving broad generalization across related tasks, new challenges emerge. One primary source of these challenges is the congruence in the underlying dynamics of different tasks, which exhibit distinct reward structures. To illustrate, two tasks might involve navigating a robot, but in one, the robot is rewarded for speed, while in the other, it is rewarded for caution. This tendency towards specialization harms the wider applicability of DRL in real-world scenarios. Developing generalizable agents that recognize and handle these subtle variations continues to be an area of keen interest and exploration in DRL.
@@
For designing such generalizable agents, a promising pathway is to train a general world model Ha and Schmidhuber (2018); Hafner et al. (2019b) that helps the agent understand the world and make the decision. In this research, we first provide insight into how a unified dynamic model can help improve the sample efficiency in handling the distribution of tasks, especially by utilizing similar structures of trajectories sampled from different tasks (Theorem 1). Consequently, we propose a novel framework named Task Aware Dreamer (TAD), which extends existing world models into the task generalization setting. As current world models are primarily designed for the single-task setting (left of Fig. 1(b)), we consider the corresponding probabilistic graphical model of the task distribution setting (right of Fig. 1(b)) and extend single-task world model into novel reward-informed world models, which can capture learnable structures in the task distribution. To optimize these models, we compute the variational lower bound of the data log-likelihood while incorporating a novel term that categorizes tasks based on their states, serving as the primary training objective for the reward-informed world models.
@@
Furthermore, we conduct detailed analyses of the components of TAD, which are simple but effective for handling task distribution. First, we point out that the reward-informed policy used in TAD is necessary for task generalization. While prior studies Ghosh et al. (2021) have highlighted that generalizing to unseen environments results in partial observability, rendering all Markovian policies suboptimal, they often overlook the relationship between the task distribution and the expressiveness of the chosen policy hypothesis set. To bridge this gap, we propose a novel metric of Task Distribution Relevance (TDR), which encapsulates the relevance of different tasks within the distribution through their optimal Q functions. Furthermore, we prove that both the commonly adopted hypothesis set $\Pi_{1}$ of Markovian policies and the set $\Pi_{2}$ of policies that encode historical states and actions, are sub-optimal under the task-distribution setting and struggle to generalize. This sub-optimality is proven to be related to TDR (Theorem 3), i.e., for task distributions with high TDR, the performance of these two policy hypotheses might degenerate significantly as they fail to identify the current task, a phenomenon we also demonstrate by experiments (Sec. 5.2). Consequently, the reward-informed policy hypothesis $\Pi_{3}$ utilized in TAD is necessary for handling task generalization. Besides the policy chosen, we also discuss the novel optimization term, i.e., categorizing tasks based on hidden state, in TAD and theoretically show that it can effectively reduce the gap between the policy return and the optimal return (Theorem 4).
@@
Extensive experiments are performed on task distributions based on DeepMind control suite Tassa et al. (2018) and MuJoCo simulator Todorov et al. (2012), which are image-based and state-based respectively. Agents are trained on specific tasks and evaluated for their generalization capabilities on unseen evaluation tasks. Our findings corroborate our analyses, indicating that $\Pi_{1}$ and $\Pi_{2}$ falter in managing task distributions characterized by a high TDR. Contrastingly, TAD excels in discerning and simultaneously managing varied tasks, significantly surpassing benchmark performance. This underscores TAD’s impressive sample efficiency and adaptability in both image-based and state-based contexts. Additionally, our ablation studies spotlight TAD’s versatility, showcasing its prowess in dynamic generalization and handling cross-embodiment tasks. In summary, our contributions are:
@@
  • We present a theoretical observation that a general world model can utilize similar structures of different tasks and improve the sample efficiency for training a generalizable agent for the task distribution.
  • Extending existing single-task world models into the task distribution setting, we then formulate a new algorithm, termed TAD, leveraging the variational lower bound for calculations and showcasing superior generalization capabilities in contrast to baselines.
  • To analyze TAD’s components, we analyze policy generalization across task distributions and introduce a new metric TDR, which quantifies the distribution relevance. Our findings indicate that Markovian policies tend to be sub-optimal, with performance gaps correlating with TDR. Further, we demonstrate that the new optimization term in TAD is effective for task generalization.
@@ 2.1 Generalization in Reinforcement Learning
Contemporary RL methodologies consistently struggle when tasked with generalizing to new environments Song et al. (2019). Prior studies have crafted an array of training methodologies, such as loss regularization Cobbe et al. (2019); Wang et al. (2020), successor representation Touati and Ollivier (2021); Touati et al. (2022), network architecture design Lee et al. (2019); Raileanu and Fergus (2021), and data augmentation Raileanu et al. (2021); Hansen and Wang (2021). Despite this, there are few investigations into the mechanism behind policy generalization in unseen environments and its relationship with the distribution of those environments. Ghosh et al. took the initial step in demonstrating that generalizing to unseen environments introduces partial observability, thereby rendering deterministic Markovian policies sub-optimal. Also, some studies Lee et al. (2020); Ghosh et al. (2021) have experimentally indicated that stochastic or non-Markovian policies can yield superior generalization compared to deterministic Markovian policies. Nevertheless, expressive capabilities of differing hypothesis sets and their connection to the environment distribution remain understudied areas. This knowledge gap is significant for understanding current algorithms and the design of more robust algorithms with enhanced generalization capacities.
@@
There are several topics closely related to generalization, like multi-task RL and meta RL. Multi-task RL Yang et al. (2020); Sodhani et al. (2021); Lee et al. (2022) primarily aims for the trained agent to excel across all training tasks. For boosting the generalization, Meta RL seeks to enable the trained agents to adapt to new tasks with as little data as possible, including gradient-based Finn et al. (2017) and context-based methods Duan et al. (2016); Rakelly et al. (2019). There are also some model-based methods Nagabandi et al. (2018); Clavera et al. (2018); Zhao et al. (2021) utilize learned models to boost the sample efficiency. Though some context-based methods like VariBAD Zintgraf et al. (2019) show zero-shot generalization ability, it is still significant to analyze the generalization in RL and design corresponding algorithms.
2.2 World Models
@@
The concept of world models Ha and Schmidhuber (2018) aims to better learn environmental representations, which has potential advantages for generalization as it can capture invariant features across tasks. Classical methods utilize the Recurrent State Space Model (RSSM) Hafner et al. (2019b) for planning like PlaNet Hafner et al. (2019b) and policy learning like Dreamer Hafner et al. (2019a, 2020, 2023). Subsequent research explored reconstruction-free world models Deng et al. (2022), preserving representational capacity by avoiding reconstructing task-irrelevant elements. Also, techniques like temporal predictive coding Nguyen et al. (2021), cooperative reconstruction Fu et al. (2021), and Denoised MDP Wang et al. (2022) have been proposed for more effective task-relevant information encoding. Recent works Chen et al. (2021); Robine et al. (2023) integrate Transformer Vaswani et al. (2017) into world models for capturing long-term dependencies. Other studies propose using world models to extract environmental invariant features by learning from videos Seo et al. (2022) or exploration Sekar et al. (2020); Xu et al. (2022), and then fine-tuning to new tasks. However, most world models are designed for the single-task setting and struggle to manage multiple tasks without fine-tuning, limiting their effectiveness for zero-shot generalization to unseen tasks.
@@ 3 Preliminary
We consider the setting with a task distribution $\mathcal{T}$ of MDPs, where different tasks own the same dynamic and different rewards. Formally, each MDP $\mathcal{M}\sim\mathcal{T}$ is represented as $\mathcal{M}=(\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R}_{\mathcal{M}},\gamma)$. Here $\mathcal{S}$ and $\mathcal{A}$ denote the state and action spaces, respectively. For $\forall(s,a)\in\mathcal{S}\times\mathcal{A}$, $\mathcal{P}(\cdot|s,a)$ is a distribution over $\mathcal{S}$, representing the dynamic, $\mathcal{R}_{\mathcal{M}}(s,a)$ is the reward function of $\mathcal{M}$, and $\gamma\in(0,1)$ is a discount factor.
Although deterministic Markovian policy is well known to be optimal for handling the single task setting, some recent work Ghosh et al. (2021) has demonstrated that generalizing to unseen environments introduces partial observability, thus deterministic Markovian policy will be sub-optimal and we need to encode more historical information. Also, previous multi-task methods Rakelly et al. (2019); Zintgraf et al. (2019) will utilize all historical information to choose the action. Formally, given policy $\pi$, at each timestep $t$, the agent will use the whole history trajectory $(s_{0},a_{0},r_{0},s_{1},a_{1},...,s_{t})$ to sample action $a_{t}$, arrive at the next state $s_{t+1}\sim\mathcal{P}(\cdot|s_{t},a_{t})$, and get the current reward $r_{t}=\mathcal{R}_{\mathcal{M}}(s_{t},a_{t})$. The performance of policy $\pi$ in $\mathcal{M}$ is defined as the expected discounted return of the trajectory $\tau$: $J_{\mathcal{M}}(\pi)=\mathbb{E}_{\tau\sim\pi}\left[R(\tau)\triangleq\sum_{t=0}^{\infty}\gamma^{t}r_{t}\right]$, where $R(\tau)$ represents the discounted return of $\tau$. And we hope to find the optimal policy over $\mathcal{T}$, i.e., $\max_{\pi}\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}\left[J_{\mathcal{M}}(\pi)\right]$.
In practice, given the task distribution $\mathcal{T}$, we sample $M$ training tasks $\{\mathcal{M}_{m}\}_{m=1}^{M}$ for optimizing the agent, i.e., maximizing $\frac{1}{M}\sum_{m=1}^{M}J_{\mathcal{M}_{m}}(\pi)$. In the testing stage, we will sample $N$ unseen testing tasks $\{\mathcal{M}_{M+n}\}_{n=1}^{N}$ to evaluate its generalization ability, i.e., evaluating $\frac{1}{N}\sum_{n=1}^{N}J_{\mathcal{M}_{M+n}}(\pi)$.
@@ 4 Task Aware Dreamer
In this section, we propose a novel framework of Task Aware Dreamer (TAD) for handling task generalization, including the reward-informed world models and the optimization. Finally, we will provide theoretical analyses of designs in TAD.
@@ 4.1 Reward-Informed World Models
Our first observation is that model-based methods, in addition to proving to improve sample efficiency in the single-task setting Janner et al. (2019); Hafner et al. (2019a), are also effective in narrowing down the hypothesis space of the optimal Q function when considering task generalization:
Theorem 1 (Informally, a detailed version and proof are in Appendix B.1).
Given $M$ tasks $\{\mathcal{M}_{m}\}_{m=1}^{M}$ and corresponding dataset $D^{m}$ sampled from $\mathcal{M}_{m}$. We can consider three hypothesis classes of optimal Q function $\mathcal{H}_{1},\mathcal{H}_{2},\mathcal{H}_{3}$ as
  • $\mathcal{H}_{1}$ includes all optimal Q functions satisfying the Bellman optimality equation for all transitions in $\mathcal{D}_{m}$.
  • For any dynamics $p_{m}$ satisfying the dynamic transition in $\mathcal{D}_{m}$, $\mathcal{H}_{2}$ includes all optimal Q functions satisfying the Bellman optimality equation for all transitions consistent with $p_{m}$.
  • For any dynamics $p$ satisfying all dynamic transition in $\{\mathcal{D}_{m}\}_{m=1}^{M}$, $\mathcal{H}_{3}$ includes all optimal Q functions satisfying the Bellman optimality equation for all transitions consistent with $p$.
We have that $\mathcal{H}_{3}\subseteq\mathcal{H}_{2}\subseteq\mathcal{H}_{1}$.
Here $\mathcal{H}_{2}\subseteq\mathcal{H}_{1}$ utilizes the generalization ability of the trained world model, which is similar to the single-task setting. Moreover, $\mathcal{H}_{3}\subseteq\mathcal{H}_{2}$ is based on similar dynamic structures across different tasks, which benefits learning the general world model. In summary, this result inherits the previous result Young et al. (2022) where the reward function is unchanged and inspires us that training a general world model can narrow down the hypothesis of possible Q function as well as benefit training the agent.
As traditional world models Hafner et al. (2019b, a) mainly designed for the single-task setting, we first analyze the probabilistic graphical model of the task-distribution setting in Fig. 1(b) and further build reward-informed world models for handling the task distribution. As shown in Fig. 1(b), the reward does not only rely on the previous states and actions but also on the current task. Thus the joint distribution can be calculated as
@@
and we choose the reward-informed inference model (analyses about the necessity of utilizing reward-informed policy in task generalization is in Sec. 4.3) to approximate state posteriors as
@@
Further, we can construct the variational lower bound of the log-likelihood on the data as
@@
It is a general form of the single task setting Hafner et al. (2019b) and the detailed derivation is in Appendix B.2. Here the first two terms in Eq.(3) are for reconstructing observations, predicting rewards, and inferring states, which are similar to the single-task setting. The last novel term is dedicated to predicting the current task from historical information, which is beneficial for improving the generalization as it encourages inferring the current task context, and we use one-hot task vectors to present $\mathcal{M}$ during training. Extending RSSM via Eq.(3), our reward-informed world models consist of:
@@
Here hidden state $h_{t}$ encodes historical states, actions, and rewards, by using gated recurrent unit (GRU) Chung et al. (2014) as $f(\cdot)=\text{GRU}(\cdot)$. Then the transition model, observation model, and reward model further predict state, observation, and reward respectively. Finally, the task model infers the current task to distinguish different tasks’ trajectories.

### v4

@@ Abstract
A long-standing goal of reinforcement learning is to acquire agents that can learn on various training tasks and generalize well on unseen tasks that may share a similar dynamic but with different reward functions. The ability to generalize across tasks is significant for real-world applications where the robot needs to adapt to varying reward mechanisms with the same embodiment. In this work, we first indicate that training general world models can utilize similar structures in these tasks and benefit training more generalizable agents. Extending world models into the task generalization setting, we introduce a novel method named Task Aware Dreamer (TAD), which integrates reward-informed features to identify consistent latent characteristics across tasks. Within TAD, we compute the variational lower bound of sample data log-likelihood, which introduces a new term designed to differentiate tasks using their states, as the optimization objective of our reward-informed world models. To demonstrate the advantages of the reward-informed policy utilized in TAD for handling the task distribution, we introduce a novel metric named Task Distribution Relevance (TDR) which quantitatively measures the relevance of different tasks. For tasks exhibiting a high TDR, i.e., these tasks differ significantly, we illustrate that Markovian policies struggle to distinguish them, thus it is necessary to utilize reward-informed policies in TAD. Extensive experiments in both image-based and state-based settings show that TAD can significantly improve the performance of handling different tasks meanwhile, especially for those with high TDR, and display a strong generalization ability to unseen tasks.
Figure 1: An overview. Given a task distribution, we train the agent in training tasks and hope it to zero-shot generalize to test tasks. For improving the generalization, we propose TAD, which utilizes $\Pi_{3}$ to encode all historical information for inferring the current task and novel reward-informed world models for capturing invariant latent features.
@@ 1 Introduction
Deep Reinforcement Learning (DRL) has demonstrated significant advancements in diverse fields (Mnih et al. 2016; Silver et al. 2016), and a key factor in these achievements is an agent’s proficiency in assimilating lessons from special training tasks. This tendency towards specialization harms the wider real-world application of DRL, where broad generalization across various tasks is necessary. One primary challenge in generalization is the correspondence in the underlying dynamics of different tasks, which exhibit distinct reward structures. For example, we may need to control the same robot to handle different tasks, which can be rewarded for speed or caution in a navigating robot. Developing generalizable agents that recognize and handle these subtle variations is still an area of keen interest and exploration in DRL.
@@
For developing such generalizable agents, a promising pathway is to train general world models (Ha and Schmidhuber 2018; Hafner et al. 2019b, a) that help the agent understand the world and make decisions. In this work, we first provide insight that general world models benefit improving the sample efficiency in handling the task distribution, especially by utilizing similar dynamic structures of trajectories sampled from different tasks (Theorem 1). Consequently, we propose a novel framework named Task Aware Dreamer (TAD) to improve the generalization ability of agents via general world models. As current world models are primarily designed for the single-task setting (left of Fig. 2), we consider the corresponding probabilistic graphical model for the task-distribution setting (right of Fig. 2) and propose novel reward-informed world models that can capture invariant structures across tasks in the distribution. Then we compute the variational lower bound of the data log-likelihood as the primary training objective for optimizing the reward-informed world models, which incorporate a novel task context term that categorizes tasks based on their states. In practice, we implement this optimization objective via two alternative methods: cross-entropy (TAD-CE) and supervised-contrastive (TAD-SC).
Figure 2: Probabilistic graphical model designs for the single-task setting (left) and the task-distribution setting (right). The latter inspires the design of reward-informed world models. Solid and dashed lines represent the generative process and the inference model, respectively.
@@
Moreover, we theoretically explain TAD’s components, like reward-informed policies and task optimization terms, are effective for handling task distribution. About the policy hypothesis, although several meta RL methods (Zintgraf et al. 2019; Rimon et al. 2024) have chosen the reward-informed policy hypothesis $\Pi_{3}$, the relationship between the task distribution and the expressiveness of the policy hypothesis set is still unclear. In other words, why the commonly adopted hypothesis set $\Pi_{1}$ of Markovian policies and the set $\Pi_{2}$ of policies encoding historical states and actions are not suitable for task distribution? To answer this question, we propose a novel metric of Task Distribution Relevance (TDR), encapsulating the relevance of different tasks within the distribution through their optimal Q functions. We then prove that both $\Pi_{1}$ and $\Pi_{2}$ are sub-optimal under the task-distribution setting. This sub-optimality is related to TDR (Theorem 3), i.e., for task distributions with high TDR, the performance of these two policy hypotheses might degenerate significantly, a phenomenon we also demonstrate in experiments (Sec. 5.2). This result explains why TAD and previous meta RL methods choose $\Pi_{3}$ for handling the task distribution. Besides the policy chosen, we also discuss our task optimization term in TAD and prove that it can effectively reduce the gap between the policy return and the optimal return (Theorem 4).
@@
We evaluate the task generalization ability of TAD in extensive experiments, including DeepMind control suite (Tassa et al. 2018) and MuJoCo (Todorov, Erez, and Tassa 2012), which are image-based and state-based respectively. Agents are trained on various tasks and evaluated on unseen tasks. Results corroborate our analyses, indicating that $\Pi_{1}$ and $\Pi_{2}$ falter in managing task distributions characterized by a high TDR. Contrastingly, TAD excels in simultaneously managing varied tasks and outperforms all baselines, including SOTA model-based meta RL method MAMBA (Rimon et al. 2024). Additionally, our ablation studies highlight TAD’s versatility, showing its prowess in dynamic generalization and handling cross-embodiment tasks. Overall, our contributions include:
@@
  • We present theoretical insights that general world models can utilize similar structures across tasks and improve the sample efficiency for task generalization (Sec. 4.1).
  • Our TAD extends world models for the task distribution with the corresponding variational lower bound (Sec. 4.2).
  • We theoretically analyze TAD’s components are effective for task generalization, with a novel metric TDR to quantify the distribution relevance (Sec. 4.3).
  • Extensive experiments show that TAD can outperform various SOTA baselines and exhibit better generalization flexibility over image-based and state-based settings (Sec. 5).
@@ 2 Related Work
Current RL methods always struggle to generalize to new tasks (Song et al. 2019). Prior works have studied an array of training strategies, like loss regularization (Cobbe et al. 2019; Wang et al. 2020), successor representation (Touati and Ollivier 2021; Touati, Rapin, and Ollivier 2022), network architecture design (Lee et al. 2019; Raileanu and Fergus 2021), lifelong learning (Chen and Liu 2018; Mendez, van Seijen, and Eaton 2022), data augmentation (Raileanu et al. 2021; Hansen and Wang 2021), etc. Besides these, some works investigate the connection between policy generalization and the distribution of those environments. Ghosh et al. demonstrates that generalizing to unseen environments introduces partial observability, thereby rendering deterministic Markovian policies sub-optimal. Also, some studies (Lee et al. 2020; Ghosh et al. 2021) experimentally indicate that stochastic or non-Markovian policies can improve the generalization ability. However, the expressive abilities of differing hypothesis sets and their connection to the environment distribution, which are significant for developing more generalizable agents, remain understudied.
Multi-task RL and Meta RL.
@@
These two topics are closely related to generalization in RL. Multi-task RL (Yang et al. 2020; Sodhani, Zhang, and Pineau 2021; Lee et al. 2022; Xu et al. 2022a) primarily aims to excel across all training tasks but is difficult to zero-shot generalize to unseen tasks. For boosting the generalization, Meta RL seeks to enable the trained agents to adapt to new tasks with few episodes, including gradient-based (Finn, Abbeel, and Levine 2017) and context-based methods (Duan et al. 2016; Rakelly et al. 2019). There are also some model-based methods (Nagabandi et al. 2018; Rimon et al. 2024) utilize learned models to boost the sample efficiency. Though some context-based methods like VariBAD (Zintgraf et al. 2019) show zero-shot generalization ability, it is still significant to directly analyze the generalization in RL and design corresponding algorithms.
World Models.
@@
World models (Ha and Schmidhuber 2018) aims to better learn environmental representations, which has potential advantages for generalization as it can capture invariant features across tasks. Classical methods utilize the Recurrent State Space Model (RSSM) (Hafner et al. 2019b) for planning (Hafner et al. 2019b) and policy learning (Hafner et al. 2019a, 2020, 2023). Subsequent research explored reconstruction-free world models (Deng, Jang, and Ahn 2022), temporal predictive coding (Nguyen et al. 2021), cooperative reconstruction (Fu et al. 2021), and Denoised MDP (Wang et al. 2022) for more effective task-relevant information encoding. World models are also utilized to extract environmental invariant features by learning from videos (Seo et al. 2022) or exploration (Sekar et al. 2020; Xu et al. 2022b), and then fine-tuning to new tasks. However, most world models are designed for the single-task setting and struggle to manage multiple tasks without fine-tuning, limiting their effectiveness for zero-shot generalization to unseen tasks.
@@ 3 Preliminary
We consider the setting with a task distribution $\mathcal{T}$ of Partially observable Markov decision processes (POMDPs), where different tasks own the same dynamic and different rewards. Formally, each POMDP $\mathcal{M}\sim\mathcal{T}$ can be represented as $\mathcal{M}=(\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R}_{\mathcal{M}},\Omega,\mathcal{O})$. Here $\mathcal{S}$ and $\mathcal{A}$ denote the state and action spaces, respectively. For $\forall(s,a)\in\mathcal{S}\times\mathcal{A}$, $\mathcal{P}(\cdot|s,a)$ is the state transition probability that is Markovian, $\mathcal{R}_{\mathcal{M}}(s,a)$ is its reward function. The state is unobserved for the agent, which can only access the observation from the observation space $\Omega$ calculated by the observation function $\mathcal{O}(\cdot|s)$.
Following previous meta RL and generalization methods, we consider policies that encoder all historical information (we prove its necessity in Sec. 4.3). Formally, at each timestep $t$, the agent with the policy $\pi$ will use the whole history trajectory $(o_{0},a_{0},r_{0},o_{1},...,o_{t})$ to sample action $a_{t}$, arrive at the next state $s_{t+1}\sim\mathcal{P}(\cdot|s_{t},a_{t})$, and get the current reward $r_{t}=\mathcal{R}_{\mathcal{M}}(s_{t},a_{t})$. The performance of policy $\pi$ in $\mathcal{M}$ is defined as the expected discounted return: $J_{\mathcal{M}}(\pi)=\mathbb{E}_{\tau\sim\pi}\left[R(\tau)\triangleq\sum_{t=0}^{\infty}\gamma^{t}r_{t}\right]$. Our objective is to maximize the return over $\mathcal{T}$, i.e., $\max_{\pi}\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}\left[J_{\mathcal{M}}(\pi)\right]$.
In practice, given the task distribution $\mathcal{T}$, we sample $M$ training tasks $\{\mathcal{M}_{m}\}_{m=1}^{M}$ for optimizing the agent, i.e., maximizing $\frac{1}{M}\sum_{m=1}^{M}J_{\mathcal{M}_{m}}(\pi)$. In the testing stage, we will sample $N$ unseen test tasks $\{\mathcal{M}_{M+n}\}_{n=1}^{N}$ to evaluate its generalization ability, i.e., evaluating $\frac{1}{N}\sum_{n=1}^{N}J_{\mathcal{M}_{M+n}}(\pi)$.
@@ 4 Task Aware Dreamer
In this section, we first demonstrate that general world models benefit task generalization. Then we introduce our reward-informed world models and propose a novel framework of Task Aware Dreamer (TAD) for handling task generalization. Finally, we provide theoretical analyses of designs in TAD.
@@ 4.1 Reward-Informed World Models
Our first observation is that general world models are effective in narrowing down the hypothesis space of the optimal Q function when handling task generalization:
Theorem 1 (Proof in Appendix A.1).
Set $\mathcal{Q}$ as the space of observation-action Q functions. Given $M$ tasks $\{\mathcal{M}_{m}\}_{m=1}^{M}$ and corresponding dataset $\mathcal{D}_{m}=\{(o_{t}^{m},a_{t}^{m},r_{t}^{m},o_{t+1}^{m})\}$, we set the product space $\mathcal{H}=\mathcal{Q}^{M}$ composed of M spaces, i.e., $\forall\{q_{m}\}_{m=1}^{M}\in\mathcal{H}$, $q_{m}:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ belongs to $\mathcal{Q}$. Considering the following three hypothesis classes $\mathcal{H}_{1},\mathcal{H}_{2},\mathcal{H}_{3}\subseteq\mathcal{H}$:
$\begin{split}\mathcal{H}_{1}=&\{(q_{m})_{m=1}^{M}|q_{m}(o_{t}^{m},a_{t}^{m})=r_{t}^{m}+\gamma\max_{a^{\prime}}q_{m}(o_{t+1}^{m},a^{\prime})\}\\ \mathcal{H}_{2}=&\{(q_{m})_{m=1}^{M}|\exists(p_{m})_{m=1}^{M}:p_{m}(o_{t}^{m},a_{t}^{m})=o_{t+1}^{m},\\ &\exists(r_{m})_{m=1}^{M}:r_{m}(o_{t}^{m},a_{t}^{m})=r_{t}^{m},\\ &q_{m}(o,a)=r_{m}(o,a)+\gamma\max_{a^{\prime}}q_{m}(p_{m}(o,a),a^{\prime}),\forall o,a\}\\ \mathcal{H}_{3}=&\{(q_{m})_{m=1}^{M}|\exists p:p(o_{t}^{m},a_{t}^{m})=o_{t+1}^{m},\\ &\exists(r_{m})_{m=1}^{M}:r_{m}(o_{t}^{m},a_{t}^{m})=r_{t}^{m},\\ &q_{m}(o,a)=r_{m}(o,a)+\gamma\max_{a^{\prime}}q_{m}(p(o,a),a^{\prime}),\forall o,a\}.\end{split}$
Here $\mathcal{H}_{1}$ own Q functions satisfying optimal Bellman equation with data from $\mathcal{D}_{m}$, while $\mathcal{H}_{2}$ and $\mathcal{H}_{3}$ own Q functions satisfying optimal Bellman function with data from world models, which are trained with the single dataset and all datasets, respectively. Consequently, $\mathcal{H}_{2}\subseteq\mathcal{H}_{1}$ utilizes the generalization ability of world models, which extends the results in the fixed-task setting (Young et al. 2023). Moreover, $\mathcal{H}_{3}\subseteq\mathcal{H}_{2}$ illustrate similar dynamic structures across different tasks benefit learning general world models. Consequently, training general world models can narrow down the hypothesis of possible Q function and benefit agent training.
As existing world models (Hafner et al. 2019b, a) are mainly designed for the single-task setting, we propose reward-informed world models for the task-distribution setting. We first analyze the probabilistic graphical model of the task-distribution setting shown in Fig. 2, where rewards do not only rely on the previous states and actions but also the current task. Thus we can calculate the joint distribution as
@@
and we choose the reward-informed inference model (the necessity of reward-informed policy in task generalization is discussed in Sec. 4.3) to approximate state posteriors as
@@
Based on this inference model, we can construct the variational lower bound of the log-likelihood on the data as
@@
This result is a general form of the single-task setting (Hafner et al. 2019b) and its derivation is in Appendix A.2. The first two terms in Eq.(LABEL:eq_elbo) are for reconstructing observations, predicting rewards, and inferring states, which are similar to the single-task setting. The last novel term is dedicated to predicting the current task from historical information, which is beneficial for improving the generalization as it encourages inferring the current task context. Extending RSSM via Eq.(LABEL:eq_elbo), our reward-informed world models consist of:
@@
Here hidden state $h_{t}$ encodes historical states, actions, and rewards, by using gated recurrent unit (GRU) (Chung et al. 2014) as $f(\cdot)=\text{GRU}(\cdot)$. Then the transition model, observation model, reward model, and task model further predict state, observation, reward, and task context respectively.